In how many different ways can 5 people arrange themselves in the 5 seats of a car for a trip if only 2 of the people can drive?
(A) 12 (B) 15 (C) 26 (D) 48 (E) 120

Answer by jim_thompson5910(35256)

2*4! = 2*24 = 48